40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 16 19:48:52 2016, Lydia, Update, General, Alignment status 
    Reply  Sat Sep 17 02:00:23 2016, rana, Update, SUS, Alignment status 
       Reply  Sat Sep 17 16:51:01 2016, Lydia, Update, SUS, Alignment status Screenshot_from_2016-09-17_16-45-00.png
          Reply  Sun Sep 18 16:18:05 2016, rana, Update, SUS, Alignment status 
             Reply  Mon Sep 19 11:11:43 2016, ericq, Update, SUS, Alignment status 
                Reply  Mon Sep 19 22:03:10 2016, ericq, Update, General, Further recovery progress 
                   Reply  Tue Sep 20 10:45:06 2016, rana, Update, General, Further recovery progress 
                      Reply  Wed Sep 21 01:08:02 2016, ericq, Update, General, Further recovery progress 
                         Reply  Wed Sep 21 09:04:57 2016, Steve, Update, General, 8 hours recovery progress 8hrs.png
                            Reply  Thu Sep 22 15:44:21 2016, Steve, Update, General, ITMX magnets are stucked again StuckAgain.png
                               Reply  Thu Sep 22 20:01:47 2016, Lydia, Update, General, ITMX freed, all optics kicked 
                                  Reply  Thu Sep 22 22:52:08 2016, ericq, Update, General, Damping found to be on 
Message ID: 12513     Entry time: Thu Sep 22 20:01:47 2016     In reply to: 12512     Reply to this: 12515
Author: Lydia 
Type: Update 
Category: General 
Subject: ITMX freed, all optics kicked 

Rana came by and freed ITMX again. I think it shouldn't be a problem for me to free it if it happens again. 

In hopes of getting better SNR on the free swing spectra, we kicked all optics at around 7pm. The damping should come back on a little after midnight. ITMX did not get stuck after this kick. 

ELOG V3.1.3-